OpenStudy (anonymous):

2x+y=7 4x+2y-14=0 I got they are parallel. Is that right?

OpenStudy (anonymous):

yup, but not only that they are the same line, if u put the equations into y=mx+b form, it is y=-2x+7 and 2y=-4x+14 AKA y=-2x+7